The following data relates to the economy of a country over a one-year period.

K’B

Subsidies ………………………………………… 1 000

Gross domestic fixed capital formation……………. 2 400

Exports of goods and services ……………………. 2 000

Government final consumption ……………………. 3 000

Property income from abroad …………………… 300

Imports of goods and services ……………………. 2 500

Value of physical decrease in stocks …………… 10

Consumer’s expenditure ……………………. 8 000

Capital consumption/Depreciation ………………… 1 500

Taxes on expenditure……………………………... 1 750

Property income paid abroad ……………………. 500

Required

Calculate the following from the above data:

(a) Gross domestic product at market prices (5 marks)

(b) Gross domestic product at factor cost (5 marks)

(c) Gross national product at factor cost (5 marks)

(d) Net national product at factor cost (5 marks)


1
Expert's answer
2021-10-31T18:25:03-0400

(a)

GDPMP = Government final consumption + private final consumption + gross domestic capital formation + net exports

"\implies 3000+8000+(2400-10)+(2000-2500)=12,890"


(b)

"GDP_{FC}=GDP_{MP} -(net\space indirect\space taxes)"

"GDP_{FC}" "=1750-1000=750"

"\implies 12890-750\\=12,140"

Net indirect taxes = indirect taxes - subsidies


(c)

Gross national product at factor cost (GNPFC)

"=GDP_{FC}+NFIA\\=12140+(300-500)\\=11940"

NFIA=Net factor income from abroad

=property income from abroad - property income paid abroad


(d)

Net national product at factor cost (NNPFC)

"\implies GNP_{FC}-Depreciation\\\implies11940-1500\\\implies10440\\=NNP_{FC}"